OpenStudy (anonymous):

2 The sum of a number and its reciprocal is 51. Find the number.

OpenStudy (tyteen4a03):

The equation you're looking for is x + (1/x) = 51.

OpenStudy (jiteshmeghwal9):

Let the number be 'x'.then, its reciprocal=\(\Large{1 \over x}\) A/q, \(\Large{\implies {{x+{1\over x}={51}}}}\) \(\Large{\implies{{x^2+1\over x}=51}}\)

OpenStudy (jiteshmeghwal9):

i think u can do this now :)

OpenStudy (anonymous):

hmm I made it to x^2-51x+1=0 but whats the next step?

OpenStudy (jiteshmeghwal9):

u can solve this quadratic equation by either quadratic formula method or factorization method if possible.

OpenStudy (jiteshmeghwal9):

so after looking @ the equation we see that it can't be done by factorization method

OpenStudy (jiteshmeghwal9):

hence, we use quadratic formula method

OpenStudy (anonymous):

Can you help me check if the answer is x = 2.6180339887499 ?

OpenStudy (tyteen4a03):

I got 50.9804. The exact form is x = 2/(51+7 sqrt(53))

OpenStudy (anonymous):

hmm what if the recipicol is 5 and 1/5?

OpenStudy (tyteen4a03):

5 + 1/5 would be 6/5.